OpenStudy (anonymous):

simplify the expression x^7*2/x^5

OpenStudy (anonymous):

is it \[x ^{7*2}/x ^{5}\] ?

OpenStudy (anonymous):

the 2/x^5 is a faction

OpenStudy (anonymous):

so it is \[x ^{7} \times 2\div x ^{5}\] ? is this correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so \[x^{7}\times2 = x \times x \times x \times x \times x \times x \times x \times 2\]

OpenStudy (anonymous):

i do not get that

OpenStudy (anonymous):

what is \[2^{2}\]

OpenStudy (anonymous):

4

OpenStudy (anonymous):

yes, so 2^2 = 2*2 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

similarly 2^3 = 2*2*2 and so on. so x^7 = x*x*x*x*x*x*x

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so x^7/x^5 = x*x*x*x*x*x*x / x*x*x*x*x = x*x = x^2

OpenStudy (anonymous):

so 2 * x^7/x^5 = 2x^2

OpenStudy (anonymous):

in general, \[x ^{a}/x ^{b} = x ^{a-b}\]

OpenStudy (anonymous):

so \[x ^{7}/x ^{5} = x ^{(7-5)} = x ^{2}\]

OpenStudy (anonymous):

got it now?

OpenStudy (anonymous):

yea a little